OpenStudy (darkbluechocobo):

The following lines are parallel to each other: 5x − 2y = -7 and -2y + 5x = 17

OpenStudy (darkbluechocobo):

is this true?

OpenStudy (briensmarandache):

parallel lines have slopes = to each other, y =(5/2)x+7/2 and y=(5/2)x-17/2 yes

OpenStudy (darkbluechocobo):

wait im confused about the 17/2 and 7/2 they arent equal though

OpenStudy (darkbluechocobo):

or are just saying that the slopes have to be equal

OpenStudy (briensmarandache):

the slopes have to be =, if the y intercepts were also = then it would just be the same line

OpenStudy (darkbluechocobo):

*facepalm* that is true thank you for explaining it

OpenStudy (briensmarandache):

no prob!!!